OpenStudy (pleasehelpmenow):

Solve the equation by using substituition and get a free medal! 2y+x=4 y-x=-7

OpenStudy (pleasehelpmenow):

please show work

OpenStudy (akshay_budhkar):

y-x=-7 x=y+7 substituting in other equation 2y+x=4 2y+y+7=4 3y=-3 y=-1 therefore x=y+7 x=-1+7 x=6 so the solution is x=6,y=-1

OpenStudy (akshay_budhkar):

you get it?